OpenStudy (anonymous):

f(x)=3x+5 and g(x)=-4x+7 find (fog)(-4)

OpenStudy (anonymous):

Given: f(x)=3x+5 and g(x)=-4x+7 Now in order to find fog(-4) first we shall find fog(x) to do this we shall substitute x= g(x)=-4x+7 in f(x) f(x)=3x+5 i.e. f(g(x))=3(-4x+7)+5 = -12x +21 +5 = -12x+26 hence f(g(x))= -12x+26 Now we substitute x=-4, we find: f(g(-4))= -12(-4)+26 = 48 +26 = 74 f(g(-4)) = 74 or fog(-4) =74 @iedoubina